# SVDQuant: Absorbing Outliers by Low-Rank Components for 4-Bit Diffusion Models

### v2

@@ 4.2 SVDQuant: Absorbing Outliers via Low-Rank Branch
Migrate outliers from activation to weight. Smoothing (Xiao et al., 2023; Lin et al., 2024a) is an effective approach for reducing outliers. We can smooth outliers in activations by scaling down the input ${\bm{X}}$ and adjusting the weight matrix ${\bm{W}}$ correspondingly using a per-channel smoothing factor ${\bm{\lambda}}\in{\mathbb{R}}^{m}$. As shown in Figure 4(a)(c), the smoothed input $\hat{{\bm{X}}}={\bm{X}}\cdot\text{diag}({\bm{\lambda}})^{-1}$ exhibits reduced magnitude and fewer outliers, resulting in lower input quantization error. However, in Figure 4(b)(d), the transformed weight $\hat{{\bm{W}}}={\bm{W}}\cdot\text{diag}({\bm{\lambda}})$ has a significant increase in both magnitude and the presence of outliers, which in turn raises the weight quantization error. Consequently, the overall error reduction is limited.
Absorb magnified weight outliers with a low-rank branch. Our core insight is to introduce a 16-bit low-rank branch and further migrate the weight quantization difficulty to this branch. Specifically, we decompose the transformed weight as $\hat{{\bm{W}}}={\bm{L}}_{1}{\bm{L}}_{2}+{\bm{R}}$, where ${\bm{L}}_{1}\in{\mathbb{R}}^{m\times r}$ and ${\bm{L}}_{2}\in{\mathbb{R}}^{r\times n}$ are two low-rank factors of rank $r$, and ${\bm{R}}$ is the residual. Then ${\bm{X}}{\bm{W}}$ can be approximated as
@@
For any tensor ${\bm{R}}$ and quantization method described in Equation 1 as $Q({\bm{R}})=s_{\bm{R}}\cdot{\bm{Q}}_{\bm{R}}$. Assuming the element of ${\bm{R}}$ follows a normal distribution, we have
$\mathbb{E}\left[\left\lVert{\bm{R}}-Q({\bm{R}})\right\rVert_{F}\right]\leq\frac{\sqrt{\log\left(\text{size}({\bm{R}})\right)\pi}}{q_{\max}}\mathbb{E}\left[\left\lVert{\bm{R}}\right\rVert_{F}\right],$ (7)
where $\text{size}({\bm{R}})$ denotes the number of elements in ${\bm{R}}$.
@@
See Appendix A.2 for the proof. From this proposition, we obtain the intuition that the quantization error $\left\lVert{\bm{R}}-Q({\bm{R}})\right\rVert_{F}$ is bounded by the magnitude of the residual $\left\lVert{\bm{R}}\right\rVert_{F}$. Thus, our goal is to find the optimal ${\bm{L}}_{1}{\bm{L}}_{2}$ that minimizes $\left\lVert{\bm{R}}\right\rVert_{F}=\left\lVert\hat{{\bm{W}}}-{\bm{L}}_{1}{\bm{L}}_{2}\right\rVert_{F}$, which can be solved by simple Singular Value Decomposition (SVD). Given the SVD of $\hat{{\bm{W}}}={\bm{U}}{\bm{\Sigma}}{\bm{V}}$, the optimal solution is ${\bm{L}}_{1}={\bm{U}}{\bm{\Sigma}}_{:,:r}$ and ${\bm{L}}_{2}={\bm{V}}_{:r,:}$. Figure 5 illustrates the singular value distribution of the original weight ${\bm{W}}$, transformed weight $\hat{{\bm{W}}}$ and residual ${\bm{R}}$. The singular values of the original weight ${\bm{W}}$ are highly imbalanced. After smoothing, the singular value distribution of $\hat{{\bm{W}}}$ becomes even sharper, with only the first several values being significantly larger. By removing these dominant values, Eckart–Young–Mirsky theorem§§§https://en.wikipedia.org/wiki/Low-rank_approximation suggests that the magnitude of the residual ${\bm{R}}$ is dramatically reduced, as $\left\lVert{\bm{R}}\right\rVert_{F}=\sqrt{\sum_{i=r+1}^{\min(m,n)}\sigma_{i}^{2}}$, compared to the original magnitude $\left\lVert\hat{{\bm{W}}}\right\rVert_{F}=\sqrt{\sum_{i=1}^{\min(m,n)}\sigma_{i}^{2}}$, where $\sigma_{i}$ is the $i$-th singular value of $\hat{{\bm{W}}}$. Furthermore, empirical observations reveal that ${\bm{R}}$ exhibits fewer outliers with a substantially compressed value range compared to $\hat{{\bm{W}}}$, as shown in Figure 4(d)(e). In practice, we can further reduce quantization errors by iteratively updating the low-rank branch through decomposing ${\bm{W}}-Q({\bm{R}})$ and adjusting ${\bm{R}}$ accordingly for several iterations, and then picking the result with the smallest error.
@@ 4.3 Nunchaku: Fusing Low-Rank and Low-Bit Branch Kernels
Although the low-rank branch introduces theoretically negligible computation, running it as a separate branch would incur significant latency overhead—approximately 50% of the 4-bit branch latency, as shown in Figure 6(a). This is because, for a small rank $r$, even though the computational cost is greatly reduced, the data sizes of input and output activations remain unchanged, shifting the bottleneck from computation to memory access. The situation deteriorates, especially when the activation cannot fit into the GPU L2 cache. For example, the up projection in the low-rank branch for QKV projection is much slower since its output exceeds the available L2 cache and results in the extra load and store operations to DRAM. Fortunately, we observe that the down projection ${\bm{L}}_{1}$ in the low-rank branch shares the same input as the quantization kernel in the low-bit branch, while the up projection ${\bm{L}}_{2}$ shares the same output as the 4-bit computation kernel, as illustrated in Figure 6(b). By fusing the down projection with the quantization kernel and the up projection with the 4-bit computation kernel, the low-rank branch can share the activations with the low-bit branch, eliminating the extra memory access and also halving the number of kernel calls. As a result, the low-rank branch adds only 5$\sim$10latency, making it nearly cost-free.
@@ 5.1 Setups
Models. We benchmark our methods using the following text-to-image models, including both the UNet (Ronneberger et al., 2015; Ho et al., 2020) and DiT (Peebles & Xie, 2023) backbones:
  • FLUX.1 (Black-Forest-Labs, 2024) is the SoTA open-sourced DiT-based diffusion model. It consists of 19 joint attention blocks (Esser et al., 2024) and 38 parallel attention blocks (Dehghani et al., 2023), totaling 12B parameters. We evaluate on both the 50-step guidance-distilled (FLUX.1-dev) and 4-step timestep-distilled (FLUX.1-schnell) variants.
  • Stable Diffusion XL (SDXL) is a widely-used UNet-based model with 2.6B parameters (Podell et al., 2024). It predicts noise with three resolution scales. The highest-resolution stage is processed entirely by ResBlocks (He et al., 2016), while the other two stages jointly use ResBlocks and attention layers. Like PixArt-$\Sigma$, SDXL employs cross-attention layers for text conditioning. We evaluate it in the 30-step setting, along with its 4-step distilled variant, SDXL-Turbo (Sauer et al., 2023).
Datasets. Following previous works (Li et al., 2023a; Zhao et al., 2024c; b), we randomly sample the prompts in COCO Captions 2024 (Chen et al., 2015) for calibration. To assess the generalization capability of our method, we adopt two distinct prompt sets with varying styles for benchmarking:
  • MJHQ-30K (Li et al., 2024a) consists of 30K samples from Midjourney with 10 common categories, 3K samples each. We uniformly select 5K prompts from this dataset to evaluate model performance on artistic image generation.
  • Densely Captioned Images (DCI) (Urbanek et al., 2024) is a dataset containing $\sim$8K images with detailed human-annotated captions, averaging over 1,000 words. For our experiments, we use the summarized version (sDCI), where captions are condensed to 77 tokens using large language models (LLMs) to accommodate diffusion models. Similarly, we randomly sample 5K prompts for efficient evaluation of realistic image generation.
@@
  • 4-bit NormalFloat (NF4) is a data type for weight-only quantization (Dettmers et al., 2023). It assumes that weights follow a normal distribution and is the information-theoretically optimal 4-bit representation. We use the community-quantized NF4 FLUX.1 models (Lllyasviel, 2024) as the baselines.
@@
Metrics. Following previous work (Li et al., 2022; 2024b), we mainly benchmark image quality and similarity to the results produced by the original 16-bit models. For the image quality assessment, we use Fréchet Inception Distance (FID, lower is better) to measure the distribution distance between the generated images and the ground-truth images (Heusel et al., 2017). Besides, we employ Image Reward (higher is better) to approximate the human rating of the generated images (Xu et al., 2024a). We use LPIPS (lower is better) to measure the perceptual similarity (Zhang et al., 2018) and Peak Signal Noise Ratio (PSNR, higher is better) to measure the numerical similarity of the images from the 16-bit models. Please refer to our Appendix B.1 for more metrics (CLIP IQA (Wang et al., 2023), CLIP Score (Hessel et al., 2021) and SSIMhttps://en.wikipedia.org/wiki/Structural_similarity_index_measure).
Table 1: Quantitative quality comparisons across different models. IR means ImageReward. Our 8-bit results closely match the quality of the 16-bit models. Moreover, our 4-bit results outperform other 4-bit baselines, effectively preserving the visual quality of 16-bit models.
@@
INT W4A4 Ours 20.1 0.898 0.394 16.2 25.1 0.922 0.434 14.9
FP W4A4 Ours 18.3 0.946 0.326 17.4 23.7 0.978 0.357 16.1
@@
INT W4A4 Ours 20.7 0.609 0.298 20.6 26.3 0.494 0.314 20.4
FP W4A4 Ours 19.0 0.607 0.294 21.0 25.4 0.480 0.312 20.7
Implementation details. For the 8-bit setting, we use per-token dynamic activation quantization and per-channel weight quantization with a low-rank branch of rank 16. For the 4-bit setting, we adopt per-group symmetric quantization for both activations and weights, along with a low-rank branch of rank 32. INT4 quantization uses a group size of 64 with 16-bit scales. FP4 quantization uses a group size of 32 with FP8 scales (Rouhani et al., 2023). For FLUX.1 models, the inputs of linear layers in adaptive normalization are kept in 16 bits (i.e., W4A16). For other models, key and value projections in the cross-attention are retained at 16 bits since their latency only covers less than 5% of total runtime.
@@ 5.2 Results
Quality results. We report the quantitative quality results in Table 1 across various models and precision levels, and show some corresponding 4-bit qualitative comparisons in Figure 7. Among all models, our 8-bit results can perfectly mirror the 16-bit results, achieving PSNR higher than 21, beating all other 8-bit baselines. On FLUX.1-dev, our INT8 PSNR even reaches 27 on MJHQ.
For 4-bit quantization, on FLUX.1, our SVDQuant surpasses the NF4 W4A16 baseline regarding Image Reward. On the schnell variant, our Image Reward even exceeds that of the original BF16 model, suggesting stronger human preference. On PixArt-$\Sigma$, while our INT4 Image Reward shows slight degradation, our FP4 model achieves an even higher score than the FP16 model. This is likely due to PixArt-$\Sigma$’s small model size (600M parameters), which is already highly compact and benefits from a smaller group size. Remarkably, both our INT4 and FP4 results consistently outperform ViDiT-Q’sOur FP16 PixArt-$\Sigma$ model is slightly different from ViDiT’s, though both offer the same quality. For fair comparisons, ViDiT-Q’s similarity results are calculated using their FP16 results. W4A8 results by a large margin across all metrics. For UNet-based models, on SDXL-Turbo, our 4-bit models significantly outperform MixDQ’s W4A8 results, and our FID scores are on par with the FP16 models, indicating no loss in performance. On SDXL, both our INT4 and FP4 results achieve comparable quality to TensorRT’s W8A8 performance, which represents the 8-bit SoTA. As shown in Figure 15 in the Appendix, our visual quality only shows minor degradation.
Figure 7: Qualitative visual results on MJHQ. Image Reward is calculated over the entire dataset. On FLUX.1 models, our 4-bit models outperform the NF4 W4A16 baselines, demonstrating superior text alignment and closer similarity to the 16-bit models. For instance, NF4 misinterprets “dinosaur style,” generating a real dinosaur. On PixArt-$\Sigma$ and SDXL-Turbo, our 4-bit results demonstrate noticeably better visual quality than ViDiT-Q’s and MixDQ’s W4A8 results.
Figure 8: SVDQuant reduces the model size of the 12B FLUX.1 by 3.6×. Additionally, our engine, Nunchaku, further cuts memory usage of the 16-bit model by 3.5× and delivers 3.0× speedups over the NF4 W4A16 baseline on both the desktop and laptop NVIDIA RTX 4090 GPUs. Remarkably, on laptop 4090, it achieves in total 10.1× speedup by eliminating CPU offloading.
Figure 9: Our INT4 model seamlessly integrates with off-the-shelf LoRAs without requiring requantization. When applying LoRAs, it matches the image quality of the original 16-bit FLUX.1-dev. See Appendix C for the text prompts.
@@
Memory save & speedup. In Figure 8, we report measured model size, memory savings, and speedup for FLUX.1. Our INT4 quantization reduces the original transformer size from 22.2 GiB to 6.1 GiB, including a 0.3 GiB overhead due to the low-rank branch, resulting in an overall 3.6× reduction. Since both weights and activations are quantized, compared to the NF4 weight-only-quantized variant, our inference engine Nunchaku even saves more memory footprint, and offers 3.0× speedup on both desktop- and laptop-level NVIDIA RTX 4090 GPUs. Notably, while the original BF16 model requires per-layer CPU offloading on the 16GB laptop 4090, our INT4 model fits entirely in GPU memory, resulting in a 10.1× speedup by avoiding offloading. We anticipate even greater speedups for FP4-quantized models on NVIDIA’s next-generation Blackwell GPUs, as they inherently support microscaling for group quantization without the need for specialized GPU kernels.
@@
Integrate with LoRA. Previous quantization methods require fusing the LoRA branches and re-quantizing the model when integrating LoRAs. In contrast, our Nunchaku eliminates redundant memory access, allowing adding a separate LoRA branch. In practice, we can fuse the LoRA branch into our low-rank branch by slightly increasing the rank, further enhancing efficiency. In Figure 9, we exhibit some visual examples of applying LoRAs of five different styles (Realism, Ghibsky Illustration, Anime, Children Sketch, and Yarn Art) to our INT4 FLUX.1-dev model. Our INT4 model successfully adapts to each style while preserving the image quality of the 16-bit version. For more visual examples, see Appendix B.2.
@@
Ablation study. In Figure 10, we present several ablation studies of SVDQuant on PixArt-$\Sigma$. First, both SVD-only and naïve quantization perform poorly in the 4-bit setting, resulting in a severe degradation of image quality. While applying smoothing to the quantization slightly improves image quality compared to naïve quantization, the overall results remain unsatisfactory. LoRC (Yao et al., 2023) introduces a low-rank branch to compensate for quantization errors, but this approach is suboptimal. Quantization errors exhibit a smooth singular value distribution. Consequently, low-rank compensation fails to effectively mitigate these errors, as discussed in Section 4.2. In contrast, we first decompose the weights and quantize only the residual. As demonstrated in Figure 5, the first several singular values are significantly larger than the rest, allowing us to shift them to the low-rank branch, which effectively reduces weight magnitude. Finally, smoothing consolidate the outliers, further enabling the low-rank branch to absorb outliers from the activations and substantially improving image quality.
Trade-off of increasing rank. Figure 11 presents the results of different rank $r$ in SVDQuant on PixArt-$\Sigma$. Increasing the rank from 16 to 64 significantly enhances image quality but increases parameter and latency overhead. In our experiments, we select a rank of 32, which offers a decent quality with minor overhead.
6 Conclusion & Discussion

### v3

@@ 4.2 SVDQuant: Absorbing Outliers via Low-Rank Branch
Migrate outliers from activation to weight. Smoothing (Xiao et al., 2023; Lin et al., 2024) is an effective approach for reducing outliers. We can smooth outliers in activations by scaling down the input ${\bm{X}}$ and adjusting the weight matrix ${\bm{W}}$ correspondingly using a per-channel smoothing factor ${\bm{\lambda}}\in{\mathbb{R}}^{m}$. As shown in Figure 4(a)(c), the smoothed input $\hat{{\bm{X}}}={\bm{X}}\cdot\text{diag}({\bm{\lambda}})^{-1}$ exhibits reduced magnitude and fewer outliers, resulting in lower input quantization error. However, in Figure 4(b)(d), the transformed weight $\hat{{\bm{W}}}={\bm{W}}\cdot\text{diag}({\bm{\lambda}})$ has a significant increase in both magnitude and the presence of outliers, which in turn raises the weight quantization error. Consequently, the overall error reduction is limited.
Absorb magnified weight outliers with a low-rank branch. Our core insight is to introduce a 16-bit low-rank branch to further migrate the weight quantization difficulty. Specifically, we decompose the transformed weight as $\hat{{\bm{W}}}={\bm{L}}_{1}{\bm{L}}_{2}+{\bm{R}}$, where ${\bm{L}}_{1}\in{\mathbb{R}}^{m\times r}$ and ${\bm{L}}_{2}\in{\mathbb{R}}^{r\times n}$ are two low-rank factors of rank $r$, and ${\bm{R}}$ is the residual. Then ${\bm{X}}{\bm{W}}$ can be approximated as
@@
For any tensor ${\bm{R}}$ and quantization method described in Equation 1 as $Q({\bm{R}})=s_{\bm{R}}\cdot{\bm{Q}}_{\bm{R}}$. Assuming the elements of ${\bm{R}}$ follow a distribution that satisfies the following regularity condition: There exists a constant $c$ such that
$\mathbb{E}\left[\max(|{\bm{R}}|)\right]\leq c\cdot\mathbb{E}\left[\left\lVert{\bm{R}}\right\rVert_{F}\right].$ (7)
Then, we have
$\mathbb{E}\left[\left\lVert{\bm{R}}-Q({\bm{R}})\right\rVert_{F}\right]\leq\frac{c\sqrt{\text{size}({\bm{R}})}}{q_{\max}}\cdot\mathbb{E}\left[\left\lVert{\bm{R}}\right\rVert_{F}\right]$ (8)
where $\text{size}({\bm{R}})$ denotes the number of elements in ${\bm{R}}$. Especially if the elements of ${\bm{R}}$ follow a normal distribution, Equation 7 holds for $c=\sqrt{\frac{\log\left(\text{size}({\bm{R}})\right)\pi}{\text{size}({\bm{R}})}}$.
@@
See Appendix 4.2 for the proof. From this proposition, we obtain the intuition that the quantization error $\left\lVert{\bm{R}}-Q({\bm{R}})\right\rVert_{F}$ is bounded by the magnitude of the residual $\left\lVert{\bm{R}}\right\rVert_{F}$. Thus, our goal is to find the optimal ${\bm{L}}_{1}{\bm{L}}_{2}$ that minimizes $\left\lVert{\bm{R}}\right\rVert_{F}=\left\lVert\hat{{\bm{W}}}-{\bm{L}}_{1}{\bm{L}}_{2}\right\rVert_{F}$, which can be solved by Singular Value Decomposition (SVD) (Eckart & Young, 1936; Mirsky, 1960). Given the SVD of $\hat{{\bm{W}}}={\bm{U}}{\bm{\Sigma}}{\bm{V}}$, the optimal solution is ${\bm{L}}_{1}={\bm{U}}{\bm{\Sigma}}_{:,:r}$ and ${\bm{L}}_{2}={\bm{V}}_{:r,:}$. Figure 5 illustrates the singular value distribution of the original weight ${\bm{W}}$, transformed weight $\hat{{\bm{W}}}$ and residual ${\bm{R}}$. The singular values of the original weight ${\bm{W}}$ are highly imbalanced. After smoothing, the singular value distribution of $\hat{{\bm{W}}}$ becomes even sharper, with only the first several values being significantly larger. By removing these dominant values, the magnitude of the residual ${\bm{R}}$ is dramatically reduced, as $\left\lVert{\bm{R}}\right\rVert_{F}=\sqrt{\sum_{i=r+1}^{\min(m,n)}\sigma_{i}^{2}}$, compared to the original magnitude $\left\lVert\hat{{\bm{W}}}\right\rVert_{F}=\sqrt{\sum_{i=1}^{\min(m,n)}\sigma_{i}^{2}}$, where $\sigma_{i}$ is the $i$-th singular value of $\hat{{\bm{W}}}$. Furthermore, Figure 4(d)(e) show that ${\bm{R}}$ exhibits fewer outliers with a substantially compressed value range compared to $\hat{{\bm{W}}}$. In practice, we further reduce quantization errors by iteratively updating the low-rank branch through decomposing ${\bm{W}}-Q({\bm{R}})$ and adjusting ${\bm{R}}$ accordingly for several iterations, and then picking the result with the smallest error.
@@ 4.3 Nunchaku: Fusing Low-Rank and Low-Bit Branch Kernels
Although the low-rank branch introduces negligible computation in theory, running it as a separate branch incurs large latency overhead—approximately 50% of the 4-bit branch latency, as shown in Figure 6(a). This occurs because, for a small rank $r$, even though the computational cost decreases significantly, the input and output activation sizes remain unchanged, shifting the bottleneck from computation to memory access. This issue worsens, especially when the activation cannot fit into the GPU L2 cache. For example, in the diffusion transformer block, the up projection in the low-rank branch for QKV projection is much slower since its output exceeds the available L2 cache, resulting in the extra DRAM load and store operations. Fortunately, the down projection ${\bm{L}}_{1}$ in the low-rank branch shares the same input as the quantization kernel in the low-bit branch, while the up projection ${\bm{L}}_{2}$ shares the same output as the 4-bit computation kernel, as illustrated in Figure 6(b). By fusing the down projection with the quantization kernel and the up projection with the 4-bit computation kernel, the low-rank branch can share the activations with the low-bit branch, eliminating the extra memory access and halving the number of kernel calls. As a result, our low-rank branch adds only 5$\sim$10latency, making it nearly cost-free.
@@ 5.1 Setups
Models. We benchmark our methods using FLUX.1 (Black-Forest-Labs, 2024), PixArt-$\Sigma$ (Chen et al., 2024a), SANA (Xie et al., 2025), Stable Diffusion XL (SDXL) (Podell et al., 2024) and SDXL-Turbo (Sauer et al., 2024), including both the UNet (Ronneberger et al., 2015; Ho et al., 2020) and DiT (Peebles & Xie, 2023) backbones. See Appendix B for more details.
Datasets. Following previous works (Li et al., 2023a; Zhao et al., 2024c; b), we randomly sample the prompts in COCO Captions 2024 (Chen et al., 2015) for calibration. To evaluate the generalization capability of our method, we sample 5K prompts from the MJHQ-30K (Li et al., 2024a) and the summarized Densely Captioned Images (sDCI) (Urbanek et al., 2024) for benchmarking. See Appendix C for more details.
@@
  • 4-bit NormalFloat (NF4) is an information-theoretically optimal 4-bit data type for weight-only quantization (Dettmers et al., 2023), which assumes that weights follow a normal distribution. We use the community-quantized NF4 FLUX.1 models (Lllyasviel, 2024) as the baselines.
@@
Metrics. Following previous works (Li et al., 2022; 2024b), we evaluate image quality and image similarity with respect to the 16-bit models’ results. For image quality assessment, we use Fréchet Inception Distance (FID, lower is better) to measure the distribution distance between the generated images and the ground-truth images (Heusel et al., 2017; Parmar et al., 2022). Besides, we employ Image Reward (higher is better) to approximate the human rating of the generated images (Xu et al., 2024a). We use LPIPS (lower is better) to measure the perceptual similarity (Zhang et al., 2018) and Peak Signal Noise Ratio (PSNR, higher is better) to measure the numerical similarity of the images from the 16-bit models. Please refer to our Appendix E.1 for more metrics (CLIP IQA (Wang et al., 2023b), CLIP Score (Hessel et al., 2021) and SSIM§§§https://en.wikipedia.org/wiki/Structural_similarity_index_measure).
Table 1: Quantitative quality comparisons across different models. RTN stands for round-to-nearest. IR means ImageReward. Our 8-bit results closely match the quality of the 16-bit models. Moreover, our 4-bit results outperform other 4-bit baselines, effectively preserving the visual quality of 16-bit models.
@@
INT W4A4 Ours 19.2 0.878 0.323 17.6 25.9 0.918 0.352 16.5
NVFP W4A4 Ours 16.6 0.940 0.271 18.5 22.9 0.971 0.298 17.2
BF16 20.6 0.952 29.9 0.847
SANA INT W4A4 RTN 20.5 0.894 0.339 15.3 28.6 0.807 0.371 13.8
-1.6B INT W4A4 Ours 19.3 0.935 0.220 17.8 28.1 0.846 0.242 16.2
(20 Steps) NVFP W4A4 RTN 19.7 0.932 0.237 17.3 29.0 0.829 0.265 15.6
NVFP W4A4 Ours 20.0 0.955 0.177 19.0 29.3 0.846 0.196 17.3
@@
INT W4A4 Ours 20.6 0.601 0.288 21.0 26.2 0.477 0.307 20.7
NVFP W4A4 Ours 18.3 0.640 0.250 21.8 23.9 0.502 0.261 21.7
Implementation details. Please refer to Appendix D fore more details.
@@ 5.2 Results
Visual quality results. We report the quantitative results in Table 1 across various models and precision levels, and show some corresponding 4-bit qualitative comparisons in Figure 7. Among all models, our 8-bit results can perfectly mirror the 16-bit results, achieving PSNR higher than 21, beating all other 8-bit baselines. On FLUX.1-dev, our INT8 PSNR even reaches 27 on MJHQ.
For 4-bit quantization, NVFP4 outperforms INT4, thanks to the native hardware support of smaller microscaling group size on Blackwell. On FLUX.1, our SVDQuant consistently surpasses the NF4 W4A16 baseline regarding all metrics. For the dev variant, our method even exceeds the original BF16 model regarding Image Reward, suggesting stronger human preference. On PixArt-$\Sigma$, while our INT4 method shows slight degradation, our NVFP4 model achieves a comparable score to the FP16 model. This is likely due to PixArt-$\Sigma$’s highly compact model size (600M parameters), which benefits from a smaller group size. Remarkably, our INT4 and NVFP4 models significantly outperform ViDiT-Q’s W4A8 results by a large margin across all metrics. Note that our FP16 PixArt-$\Sigma$ model differs slightly from ViDiT’s, though both offer the same quality. For fair comparisons, ViDiT-Q’s similarity results are calculated using their FP16 results.
For UNet-based models, on SDXL-Turbo, our 4-bit models substantially outperform MixDQ W4A8, and our FID scores are on par with the FP16 models, indicating no quality loss. On SDXL, our INT4 and NVFP4 results achieve comparable quality to TensorRT’s W8A8 performance, which represents the 8-bit SoTA. As shown in Figure 14 in the Appendix, our visual quality only shows minor degradation.
Figure 7: Qualitative visual results on MJHQ. Image Reward is calculated over the entire dataset. On FLUX.1 models, our 4-bit models outperform the NF4 W4A16 baselines, demonstrating superior text alignment and closer similarity to the 16-bit models. For instance, NF4 misses the swinging chair in the top right example. On PixArt-$\Sigma$ and SDXL-Turbo, our 4-bit results demonstrate noticeably better visual quality than ViDiT-Q’s and MixDQ’s W4A8 results.
Figure 8: SVDQuant reduces the 12B FLUX.1 model size by 3.6× and cuts the 16-bit model’s memory usage by 3.5×. With Nunchaku, our INT4 model runs 3.0× faster than the NF4 W4A16 baseline on both desktop and laptop NVIDIA RTX 4090 GPUs. Notably, on the laptop 4090, it achieves a total 10.1× speedup by eliminating CPU offloading. Our NVFP4 model is also 3.1× faster than both BF16 and NF4 on the RTX 5090 GPU.
Figure 9: Our 4-bit model seamlessly integrates with off-the-shelf LoRAs without requiring requantization. When applying LoRAs, it matches the image quality of the original 16-bit FLUX.1-dev. See Appendix F for the text prompts.
@@
Memory save and speedup. In Figure 8, we report measured model size, memory savings, and speedup for FLUX.1. Our INT4 and NVFP4 quantization reduce the original transformer size from 22.2 GiB to 6.1 GiB, including a 0.3 GiB overhead due to the low-rank branch, resulting in an overall 3.6× reduction. Since both weights and activations are quantized, compared to the NF4 weight-only-quantized variant, our inference engine Nunchaku even saves more memory footprint. It offers a 3.0× speedup on both desktop- and laptop-level NVIDIA RTX 4090 GPUs with INT4 precision and a 3.1× speedup on the RTX 5090 GPU with NVFP4 precision, compared to both NF4 and the original 16-bit models. Notably, while the original BF16 model requires per-layer CPU offloading on the 16GB laptop 4090, our INT4 model fits entirely in GPU memory, resulting in a 10.1× speedup by avoiding offloading.
@@
Integrate with LoRA. Previous quantization methods require fusing the LoRA branches and re-quantizing the model when integrating LoRAs. In contrast, our Nunchaku eliminates redundant memory access, allowing adding a separate LoRA branch. In practice, we can fuse the LoRA branch into our low-rank branch by slightly increasing the rank, further enhancing efficiency. In Figure 9, we exhibit some visual examples of applying LoRAs of five different styles (Realism, Ghibsky Illustration, Anime, Children Sketch, and Yarn Art) to our INT4 FLUX.1-dev model. Our INT4 model successfully adapts to each style while preserving the image quality of the 16-bit version. For more visual examples, see Appendix E.2. For FLUX.1-schnell, we further support LoRAs from one-step conditional model pix2pix-turbo (Parmar et al., 2024), enabling additional controls like sketch. An interactive demo is available here.
@@
Ablation study. In Figure 10, we present several ablation studies of SVDQuant on PixArt-$\Sigma$. First, both SVD-only and naïve quantization perform poorly in the 4-bit setting, resulting in severe quality degradation. While applying smoothing to the quantization slightly improves image quality compared to naïve quantization, the results remain unsatisfactory. LoRC (Yao et al., 2024) introduces a low-rank branch to compensate for quantization errors, but this approach is suboptimal, as quantization errors exhibit a well-spread distribution of singular values. Consequently, low-rank compensation fails to effectively mitigate these errors, as discussed in Section 4.2. In contrast, we first decompose the weights and quantize only the residual. As demonstrated in Figure 5, the first several singular values are significantly larger than the rest, allowing us to shift them to the low-rank branch, effectively reducing weight magnitude. Finally, smoothing consolidates the outliers, enabling the low-rank branch to absorb outliers from the activations and substantially improving image quality.
Trade-off of increasing rank. Please refer to Appendix E.5 for more details.
6 Conclusion